# GmNet: Revisiting Gating Mechanisms From A Frequency View

### v2

@@ 4.1 Rethinking Current Lightweight Model Architectures From A Frequency Perspective
Figure 5: The illustration of GmNet architecture. GmNet adopts a traditional hybrid architecture, utilizing convolutional layers to down-sample the resolution and double the number of channels at each stage.
Before introducing our proposed network, we first investigate the importance of capturing high-frequency information in efficient architectures by modifying existing efficient models to incorporate Gated Linear Units (GLUs). Specifically, we select two representative architectures: the pure CNN-based MobileNetV2 sandler2018mobilenetv2 and the transformer-based EfficientFormer-V2 li2023rethinking . In both models, we replace the activation functions in their MLP blocks with a simple GLU. Detailed architectural modifications are provided in the supplementary material.
As shown in Fig. 4, we present the testing accuracy curves under varying frequency thresholds. Our results demonstrate that integrating GLUs improves classification accuracy on high-frequency components across both models, under different threshold settings. Notably, this improvement in high-frequency classification also correlates with a gain in overall performance. Furthermore, we observe that using GELU as the activation function within the GLUs enhances performance on low-frequency components, though it has a relatively minor effect on overall accuracy. These findings suggest that effectively modeling high-frequency information is more crucial for improving the performance of lightweight neural networks. It underscores the critical role of frequency-aware design in lightweight network architectures.
@@ 4.2 Architecture of GmNet
To address the limitation of low-frequency bias for current lightweight network designs, our proposed method named as GmNet integrates a simple gated linear unit into the block as illustrated in Fig. 5. GmNet offers both theoretical and practical advantages on encouraging the model to learn from a broader range of frequency regions, especially the high-frequency domain. We incorporate two depth-wise convolution layers with kernel sizes of $7\times 7$ at the beginning and end of the block respectively to facilitate the integration of low- and high-frequency information. At the core of the block, we have two $1\times 1$ convolution layers and a simple gated linear unit. We use the ReLU6 as the activation function. We vary the block numbers, input embedding channel numbers and channel expansion factors ‘ratio’ to build different sizes of GmNet, as detailed in Table 5.1.
@@ 5.1 Results in Image Classification
Implementation details. We perform image classification experiments on the ImageNet-1K dataset, adopting a standard input resolution of $224\times 224$ for both training and evaluation. All model variants are trained from scratch for $300$ epochs using the AdamW optimizer, starting with an initial learning rate of $3\times 10^{-3}$ and a batch size of $2048$. The supplementary materials provide a comprehensive overview of the training setup. For performance assessment, we convert our PyTorch models into the ONNX format to measure latency on a Mobile device (iPhone 14) and a GPU (A100). Additionally, we deploy the models on the mobile device via CoreML-Tools to further evaluate latency. Importantly, our training approach does not incorporate advanced techniques such as re-parameterization or knowledge distillation. The results presented in Table 5.1 correspond to models trained without these enhancements.
Compared with the state-of-the-art. The experimental results are presented in Table 5.1. Without any strong training strategy, GmNet delivers impressive performance compared to many state-of-the-art lightweight models. With a comparable latency on GPU, GmNet-S1 outperforms MobileV2-1.0 by $3.5\%$. Notably, GmNet-S2 achieves $78.3\%$ with only $1.9$ms on the A100 which is a remarkable achievements for the models under 1G FLOPS. GmNet-S3 outperforms RepViT-M1.0 and StarNet-S4 by $1.9\%$ and $0.9\%$ in top-1 accuracy with $1.1$ ms and $1.4$ ms faster on the GPU latency, respectively. The improvements on the speed are over $30\%$. Additionally, with similar latency, GmNet-S3 delivers a $1.7\%$ improvement on the accuracy over MobileOne-S4. GmNet-S4 achieves 2x faster compared to RepViT-M1.5 on the GPU and it surpasses MobileOne-S4 of $2.1\%$ under the similar latencies of both GPU and Mobile. LeViT-256 graham2021levit matches the accuracy of GmNet-S4 but runs twice as slow on a GPU and 16 times slower on an iPhone 14 The strong performance of GmNet can be largely attributed to the clear insights of gating mechanisms and simplest architectures. Fig. 6 further illustrates the latency-accuracy trade-off across different models. More comparisons and results of downstream tasks including objective detection and segmentation can be found in supplementary.
Variant $C_{1}$ depth ratio Params FLOPs
GmNet-S1 $40$ $[2,2,10,2]$ $[3,3,3,2]$ 3.7M 0.6G
GmNet-S2 48 $[2,2,8,3]$ $[3,3,3,2]$ 6.2M 0.9G
GmNet-S3 $48$ $[3,3,8,3]$ $[4,4,4,4]$ 7.8M 1.2G
GmNet-S4 $68$ $[3,3,11,3]$ $[4,4,4,4]$ 17.0M 2.7G
Figure 6: Trade-off between Top-1 accuracy and latency on A100. GmNet variants achieve substantially lower latency compared to related works, while maintaining competitive or superior Top-1 accuracy. This demonstrates the efficiency of our design in balancing performance and computational cost. A more extensive comparison is provided in the supplementary material.
Table 2: Comparison of Efficient Models on ImageNet-1k. Latency is evaluated across various platforms (A100 GPU and iPhone 14).
@@
FasterNet-T0 chen2023run 71.9 3.9 0.3 2.5 0.7
MobileV2-1.0 sandler2018mobilenetv2 72.0 3.4 0.3 1.7 0.9
ShuffleV2-1.5 ma2018shufflenet 72.6 3.5 0.3 2.2 1.3
EfficientFormerV2-S0 li2023rethinking 73.7 3.5 0.4 2.0 0.9
MobileNetv4-Conv-S qin2024mobilenetv4 73.8 3.8 0.2 2.2 0.9
StarNet-S2 ma2024rewrite 74.8 3.7 0.5 1.9 0.9
LSNet-T wang2025lsnet 74.9 11.4 0.3 2.9 1.8
GmNet-S1 75.5 3.7 0.6 1.6 1.0
EfficientMod-xxs ma2024efficient 76.0 4.7 0.6 2.3 18.2
Fasternet-T1 chen2023run 76.2 7.6 0.9 2.5 1.0
EfficientFormer-L1 li2022efficientformer 77.2 12.3 1.3 12.1 1.4
StarNet-S3 ma2024rewrite 77.3 5.8 0.7 2.3 1.1
MobileOne-S2 vasu2023mobileone 77.4 7.8 1.3 1.9 1.0
RepViT-M0.9 wang2024repvit 77.4 5.1 0.8 3.0 1.1
EfficientFormerV2-S1 li2023rethinking 77.9 4.5 0.7 3.4 1.1
GmNet-S2 78.3 6.2 0.9 1.9 1.1
EfficientMod-xs ma2024efficient 78.3 6.6 0.8 2.9 22.7
StarNet-S4 ma2024rewrite 78.4 7.5 1.1 3.3 1.1
SwiftFormer-S shaker2023swiftformer 78.5 6.1 1.0 3.8 1.1
RepViT-M1.0 wang2024repvit 78.6 6.8 1.2 3.6 1.1
UniRepLKNet-F ding2024unireplknet 78.6 6.2 0.9 3.1 3.5
GmNet-S3 79.3 7.8 1.2 2.1 1.3
RepViT-M1.1 wang2024repvit 79.4 8.3 1.3 5.1 1.2
MobileOne-S4 vasu2023mobileone 79.4 14.8 2.9 2.9 1.8
FastViT-S12 vasu2023fastvit 79.8 8.8 1.8 5.3 1.6
MobileNetv4-Conv-M qin2024mobilenetv4 79.9 9.2 1.0 9.2 1.4
LSNet-B wang2025lsnet 80.3 23.2 1.3 6.2 3.6
EfficientFormerV2-S2 li2023rethinking 80.4 12.7 1.3 5.4 1.6
EfficientMod-s ma2024efficient 81.0 12.9 1.4 4.5 35.3
RepViT-M1.5 wang2024repvit 81.2 14.0 2.3 6.4 1.7
LeViT-256 graham2021levit 81.5 18.9 1.1 6.7 31.4
GmNet-S4 81.5 17.0 2.7 2.9 1.9
@@ 5.2 Ablation Studies
Comparison with existing methods. As addressed in Table 3, a model should achieves strong performance across different frequency components to deliver a better overall performance.
Table 4: Comparison with recent methods. We test the models on the high-/low-frequency components with different radii on the ImageNet-1k. The highest values of each columns are highlighted.
@@
MobileOne-S2 vasu2023mobileone 77.4 35.0 11.6 6.5 36.9 2.4 53.5
EfficientMod-xs ma2024efficient 78.3 45.4 12.9 9.4 40.6 3.5 54.6
StarNet-S4 ma2024rewrite 78.4 43.3 13.8 9.4 41.3 3.4 54.8
@@
However, both pure convolutional architectures and transformers exhibit a low-frequency bias, as discussed in bai2022improving ; tang2022defects . Therefore, enhancing the performance of a lightweight model depends on its ability to more effectively capture high-frequency information. To address the advantages of GmNet on overcoming the low-frequency bias, we test some existing models on different frequency components of different radii. We select three kinds of typical lightweight methods for comparison including pure conv-based model MobileOne-S2 vasu2023mobileone , attention-based model EfficientMod-xs ma2024efficient and model also employing GLUs-like structure StarNet-S4 ma2024rewrite . As shown in Table 4, accuracies of low-frequency components are close among different models considering the overall performance. However, it shows that GmNet-S3 clearly surpass the other models in high frequency components. For example, GmNet-S3 has a $6.3\%$ improvement compared to EfficientMod-xs when $r=12$ and $2.7\%$ increase when $r=24$. For StarNet, which also uses a GLU-like structure with dual-channel FC, it struggles to effectively emphasize high-frequency signals. The simplest GLUs design can achieve a better balance between the efficiency and the effectiveness.
@@
Study on designs of the GLU. In GmNet, the gated linear unit adopts the simplest design, which can be defined as $\sigma(x)\cdot x$. For comparison, we modify the GLU design and conduct experiments to test performance on raw data as well as on decompositions at different frequency levels. As shown in the Table 5, the simplest design achieve the best performance both on effectiveness and efficiency for the overall performance. For the decomposed frequency components, we observe clear differences among various GLU designs. The GLU of $\sigma(x)\cdot x$ demonstrates significantly higher accuracy in classifying high-frequency components. For example, for $r=12$ and $r=24$, the GLU with $\sigma(x)\cdot x$ shows an improvement of $4.1$ over the LN design and $2.5$ over the DW design. This indicates that the simplest GLU design is already effective at introducing reliable high-frequency components to enhance the model’s ability to learn them. Designs aimed at smoothing information show a notable improvement in some low-frequency components. For instance, with similar overall performance, the GLUs using $\sigma(x)\cdot\text{DW}(x)$ and $\sigma(x)\cdot\text{LN}(x)$ achieve better results on low-frequency components when the radii are set to $24$, $36$, and $48$. And depth-wise convolution is more effective than layer normalization in encouraging neural networks to learn from low-frequency components which is also more efficient. For the design with the average pooling, it does not perform better in classifying high-frequency signals. This may be because $x-\text{pool}(x)$ acts as an overly aggressive high-pass filter, which does not retain the original high-frequency signals in $x$ well and instead introduces more high-frequency noise.
@@
Bandwidths analysis of convolution kernels. As discussed in the tang2022defects , the convolution layer may play roles of ’smoothing’ the feature which means it has a low-frequency bias. Experiments on studying weights of the convolution layer is insightful to give more evidences of how GLUs effect the learning of different frequency components wang2020high ; tang2022defects ; bai2022improving . In this paper, we propose using the bandwidths of convolution kernels to represent their ability of responding to different frequency components. Specifically, a wider bandwidth indicates that the kernel can process a broader range of frequencies, allowing it to capture diverse frequency components simultaneously and thereby preserve rich information from the feature. As illustrated in Figure 7, the distributions of the ReLU model suggest that its convolution kernels tend to focus on a narrow range of frequency components leading to relatively lower bandwidths. It indirectly reflects an overemphasis on high-frequency components. Although the model using GELU exhibits a better distribution in the top convolutional layers, it still has a low-frequency bias, leading to a distribution shift in the bottom convolutional layers. Compared to other activations, the enhanced bandwidth distribution of the model using ReLU6 demonstrates better generalization for this task. The properties of the convolution kernels align with the results in Table 3.
@@
Figure 7: The histogram illustrates the distribution of bandwidths of convolution kernels. We use weights of the convolution layer which under the GLU in the first block (left) and the last block (right) of the GmNet-S3. All modals are trained on the raw data of the ImageNet-1k. In general, the further the distribution shifts to the right, the stronger the convolutional kernel’s ability to capture signals of different frequencies.
6 Limitations
Discussions. Different datasets and training strategies may lead to different dynamics which has been addressed in wang2020high ; tang2022defects . For the small dataset which is used in Sec. 3, the convergence is faster which means NNs can easily capture some certain frequency components and then learning from other frequency components becomes very slow. Moreover, from the performance on very high-frequency components (for $r>36$), it is evident that neural networks face challenges in directly classifying these components, which are nearly imperceptible to the human eye. However, NNs can be sensitive to the perturbation of high-frequency noise yin2019fourier . This is also an interesting topic of revealing the robustness of NNs, but it is beyond the scope of this article.
Limitations. This work investigates that enhancing high-frequency learning is crucial for improving overall model performance for efficient models. However, this does not imply that GLUs are a universally optimal solution for all architectures. Larger models, with higher embedding dimensions and deeper structures, may inherently mitigate low-frequency bias, reducing the relative impact of such gating mechanisms. In this paper, we focus specifically on efficient networks. Additionally, in architectures with more complex structures, simply replacing the activation function with a GLU may not significantly enhance the model’s ability to capture high-frequency information.

### v3

@@ 4.1 Rethinking Current Lightweight Model Architectures From A Frequency Perspective
Figure 5: GmNet architecture. GmNet adopts a traditional hybrid architecture, utilizing convolutional layers to down-sample the resolution and double the number of channels at each stage.
Before introducing our proposed network, we first investigate the importance of capturing high-frequency information in efficient architectures by modifying existing efficient models to incorporate Gated Linear Units (GLUs). Specifically, we select one representative architecture: the pure CNN-based MobileNetV2 Sandler et al. (2018). We replace the activation functions in their MLP blocks with a simple GLU. Detailed architectural modifications are provided in the appendix.
As shown in Fig. 4, we present the testing accuracy curves under the frequency threshold $r=10$. Our results demonstrate that integrating GLUs improves classification accuracy on high-frequency components. Notably, this improvement in high-frequency classification also correlates with a gain in overall performance. Furthermore, we observe that using GELU as the activation function within the GLUs enhances performance on low-frequency components, though it has a relatively minor effect on overall accuracy. These findings suggest that effectively modeling high-frequency information is more crucial for improving the performance of lightweight neural networks. It underscores the critical role of frequency-aware design in the lightweight networks. Moreover, we also conduct similar experiments on the transformer-based model EfficientFormer-V2 Li et al. (2023) which can be found in the appendix.
@@ 4.2 Architecture of GmNet
To address the limitation of low-frequency bias for current lightweight network designs, our proposed method named as GmNet integrates a simple gated linear unit into the block as illustrated in Fig. 5. GmNet offers both conceptual and practical advantages on encouraging the model to learn from a broader range of frequency regions, especially the high-frequency domain.
GmNets employ an extremely streamlined model architecture, carefully designed to minimize both parameter count and computational speed, making them particularly suitable for deployment in resource-constrained environments. We incorporate two depth-wise convolution layers with kernel sizes of $7\times 7$ at the beginning and end of the block respectively to facilitate the integration of low- and high-frequency information. At the core of the block, we have two $1\times 1$ convolution layers and a simple gated linear unit. We use the ReLU6 as the activation function.
@@ 5.1 Results in Image Classification
Experimental Setup.
We evaluate our models on the ImageNet-1K benchmark using
Table 1: Comparison of Efficient Models on ImageNet-1k. Latency is evaluated across various platforms.
@@
FasterNet-T0 Chen et al. (2023) 71.9 3.9 0.3 2.5 0.7
MobileV2-1.0 Sandler et al. (2018) 72.0 3.4 0.3 1.7 0.9
ShuffleV2-1.5 Ma et al. (2018) 72.6 3.5 0.3 2.2 1.3
EfficientFormerV2-S0 Li et al. (2023) 73.7 3.5 0.4 2.0 0.9
MobileNetv4-Conv-S Qin et al. (2024) 73.8 3.8 0.2 2.2 0.9
StarNet-S2 Ma et al. (2024a) 74.8 3.7 0.5 1.9 0.9
LSNet-T Wang et al. (2025) 74.9 11.4 0.3 2.9 1.8
[…]
EfficientMod-xxs Ma et al. (2024b) 76.0 4.7 0.6 2.3 18.2
Fasternet-T1 Chen et al. (2023) 76.2 7.6 0.9 2.5 1.0
EfficientFormer-L1 Li et al. (2022) 77.2 12.3 1.3 12.1 1.4
StarNet-S3 Ma et al. (2024a) 77.3 5.8 0.7 2.3 1.1
MobileOne-S2 Vasu et al. (2023b) 77.4 7.8 1.3 1.9 1.0
RepViT-M0.9 Wang et al. (2024) 77.4 5.1 0.8 3.0 1.1
EfficientFormerV2-S1 Li et al. (2023) 77.9 4.5 0.7 3.4 1.1
[…]
EfficientMod-xs Ma et al. (2024b) 78.3 6.6 0.8 2.9 22.7
StarNet-S4 Ma et al. (2024a) 78.4 7.5 1.1 3.3 1.1
SwiftFormer-S Shaker et al. (2023) 78.5 6.1 1.0 3.8 1.1
RepViT-M1.0 Wang et al. (2024) 78.6 6.8 1.2 3.6 1.1
UniRepLKNet-F Ding et al. (2024) 78.6 6.2 0.9 3.1 3.5
[…]
RepViT-M1.1 Wang et al. (2024) 79.4 8.3 1.3 5.1 1.2
MobileOne-S4 Vasu et al. (2023b) 79.4 14.8 2.9 2.9 1.8
FastViT-S12 Vasu et al. (2023a) 79.8 8.8 1.8 5.3 1.6
MobileNetv4-Conv-M Qin et al. (2024) 79.9 9.2 1.0 9.2 1.4
LSNet-B Wang et al. (2025) 80.3 23.2 1.3 6.2 3.6
EfficientFormerV2-S2 Li et al. (2023) 80.4 12.7 1.3 5.4 1.6
EfficientMod-s Ma et al. (2024b) 81.0 12.9 1.4 4.5 35.3
RepViT-M1.5 Wang et al. (2024) 81.2 14.0 2.3 6.4 1.7
LeViT-256 Graham et al. (2021) 81.5 18.9 1.1 6.7 31.4
[…]
an input resolution of $224\times 224$ for both training and inference. To construct different GmNet variants, we vary the stage depth, embedding width, and expansion ratio while preserving the overall architectural design. Detailed configurations of each model scale are provided in the appendix. All networks are optimized from random initialization for $300$ epochs using the AdamW optimizer with an initial learning rate of $3\times 10^{-3}$. The global batch size is set to $2048$. Additional training hyperparameters, including data augmentation and regularization settings, are summarized in the supplementary material. For latency evaluation, we export the trained PyTorch models to ONNX format and measure inference time on both an NVIDIA A100 GPU and a mobile device (iPhone 14). To ensure fair deployment comparison, the mobile benchmarks are conducted via CoreML conversion. Importantly, no re-parameterization, distillation, or architecture search techniques are applied; all reported results correspond to standard end-to-end training.
Compared with the state-of-the-art. The experimental results are presented in Table 1. Without any strong training strategy, GmNet delivers impressive performance compared to many state-of-the-art lightweight models. With a comparable latency on GPU, GmNet-S1 outperforms MobileV2-1.0 by $3.5\%$. Notably, GmNet-S2 achieves $78.3\%$ with only 1.91.9ms on the A100 which is a remarkable achievements for the models under 1G FLOPS. GmNet-S3 outperforms RepViT-M1.0 and StarNet-S4 by $1.9\%$ and $0.9\%$ in top-1 accuracy with $1.1$ ms and
Figure 6: Trade-off between Top-1 accuracy and latency on A100.
$1.4$ ms faster on the GPU latency, respectively. The improvements on the speed are over $30\%$. Additionally, with similar latency, GmNet-S3 delivers a $1.7\%$ improvement on the accuracy over MobileOne-S4. GmNet-S4 achieves 2x faster compared to RepViT-M1.5 on the GPU and it surpasses MobileOne-S4 of $2.1\%$ under the similar latencies of both GPU and Mobile. LeViT-256 Graham et al. (2021) matches the accuracy of GmNet-S4 but runs twice as slow on a GPU and 16 times slower on an iPhone 14 The strong performance of GmNet can be largely attributed to the clear insights of gating mechanisms and simplest architectures. Fig. 6 further illustrates the latency-accuracy trade-off across different models. GmNet variants achieve substantially lower latency compared to related works, while maintaining competitive or superior Top-1 accuracy. More comparisons and results can be found in supplementary.
@@ 5.2 Ablation Studies
Comparison with existing methods from the frequency perspective. As addressed in Table 2, a model should achieves strong performance across different frequency components to deliver a better overall performance. However, both pure convolutional architectures and transformers exhibit a low-frequency bias, as discussed in Bai et al. (2022); Tang et al. (2022). Therefore, enhancing the performance of a lightweight model depends on its ability to more effectively capture high-frequency information. To address the advantages of GmNet on overcoming the low-frequency bias, we test some existing models on different frequency components of different
Table 3: Comparison with recent methods. We test models on the high-/low-frequency components on the ImageNet-1k. The highest values of each columns are highlighted.
@@
MobileOne-S2 Vasu et al. (2023b) 77.4 35.0 11.6 6.5 36.9 2.4 53.5
EfficientMod-xs Ma et al. (2024b) 78.3 45.4 12.9 9.4 40.6 3.5 54.6
StarNet-S4 Ma et al. (2024a) 78.4 43.3 13.8 9.4 41.3 3.4 54.8
@@
radii. We select three kinds of typical lightweight methods for comparison including pure conv-based model MobileOne-S2 Vasu et al. (2023b), attention-based model EfficientMod-xs Ma et al. (2024b) and model also employing GLUs-like structure StarNet-S4 Ma et al. (2024a). As shown in Table 3, accuracies of low-frequency components are close among different models considering the overall performance. However, it shows that GmNet-S3 clearly surpass the other models in high frequency components. For example, GmNet-S3 has a $6.3\%$ improvement compared to EfficientMod-xs when $r=12$ and $2.7\%$ increase when $r=24$. For StarNet, which also uses a GLU-like structure with dual-channel FC, it struggles to effectively emphasize high-frequency signals. The simplest GLUs design can achieve a better balance between the efficiency and the effectiveness.
@@
Study on designs of the GLU. In GmNet, the gated linear unit adopts the simplest design, which can be defined as $\sigma(x)\cdot x$. For comparison, we modify the GLU design and conduct experiments to test performance on raw data as well as on decompositions at different frequency levels. As shown in the Table 4, the simplest design achieve the best performance both on effectiveness and efficiency for the overall performance. For the decomposed frequency components, we observe clear differences among various GLU designs. The GLU of $\sigma(x)\cdot x$ demonstrates significantly higher accuracy in classifying high-frequency components. For example, for $r=12$ and $r=24$, the GLU with $\sigma(x)\cdot x$ shows an improvement of $4.1$ over the LN design and $2.5$ over the DW design. This indicates that the simplest GLU design is already effective at introducing reliable high-frequency components to enhance the model’s ability to learn them. Designs aimed at smoothing information show a notable improvement in some low-frequency components. For instance, with similar overall performance, the GLUs using $\sigma(x)\cdot\text{DW}(x)$ and $\sigma(x)\cdot\text{LN}(x)$ achieve better results on low-frequency components when the radii are set to $24$, $36$, and $48$. The model using a linear layer in GLUs offers performance comparable to GmNet-S3 and is adept at learning low-frequency features. However, its placement at a high-dimensional stage is problematic. This design choice leads to an excessive number of parameters and a significant increase in latency. Moreover, depth-wise convolution is more effective than layer normalization in encouraging neural networks to learn from low-frequency components which is also more efficient. For the design with the average pooling, it does not perform better in classifying high-frequency signals. This may be because $x-\text{pool}(x)$ acts as an overly aggressive high-pass filter, which does not retain the original high-frequency signals in $x$ well and instead introduces more high-frequency noise.
@@
Bandwidths analysis of convolution kernels. As discussed in the Tang et al. (2022), the convolution layer may play roles of ’smoothing’ the feature which means it has a low-frequency bias. Experiments on studying weights of the convolution layer is insightful to give more evidences of how GLUs effect the learning of different frequency components Wang et al. (2020); Tang et al. (2022); Bai et al. (2022). In this paper, we propose using the bandwidths of convolution kernels to represent their ability of responding to different frequency components. Specifically, a wider bandwidth indicates that the kernel can process a broader range of frequencies, allowing it to capture diverse frequency components simultaneously and thereby preserve rich information from the feature. As illustrated in Figure 7, the distributions of the ReLU model suggest that its convolution kernels tend to focus on a narrow range of frequency components leading to relatively lower bandwidths. It indirectly reflects an overemphasis on high-frequency components. Although the model using GELU exhibits a better distribution in the top convolutional layers, it still has a low-frequency bias, leading to a distribution shift in the bottom convolutional layers. Compared to other activations, the enhanced bandwidth distribution of the model using ReLU6 demonstrates better generalization for this task. The properties of the convolution kernels align with results in Table 2.
@@
Figure 7: The histogram illustrates the distribution of bandwidths of convolution kernels. Bandwidths represents the capability of a convolution kernel for capturing various frequency information. We use weights of the convolution layer which under the GLU in the first block (left) and the last block (right) of the GmNet-S3. All modals are trained on the raw data of the ImageNet-1k. In general, the further the distribution shifts to the right, the stronger the convolutional kernel’s ability to capture signals of different frequencies.
Contribution breakdown. We show the ablation study of different block designs and adjust the
Table 5: Contribution breakdown under matched FLOPs and parameters (mean $\pm$ std).
Variant Params (M) FLOPs (G) Top-1 Acc (%)
Baseline 7.82 1.24 71.5 $\pm$ 0.2
+ $\times7$ DWConv 7.82 1.28 78.1 $\pm$ 0.1
+ Gate (Identity) 7.82 1.24 69.2 $\pm$ 0.3
+ Gate (ReLU) 7.82 1.24 78.0 $\pm$ 0.2
+ Gate (GELU) 7.82 1.24 77.9 $\pm$ 0.1
+ Gate (ReLU6) 7.82 1.24 78.5 $\pm$ 0.1
+ ReLU6 7.82 1.24 77.9 $\pm$ 0.1
Full GmNet 7.82 1.24 79.2 $\pm$ 0.1
dimensions/number of blocks under strictly matched FLOPs, parameter count, and training settings, and report the mean ± std over three random seeds. We replace the 7×7 dwconv with a linear layer and replace the GLU with a ReLU as the baseline block. The results are shown in Table 5. Overall, Table demonstrates that the performance gains introduced by gating mechanism are distinct improvements beyond what other components alone can offer.